OpenStudy (anonymous):

Evaluate the following expression. 6-3

OpenStudy (lolacole12):

6^(-3) = ( \frac{1}{6} )^3 = \frac{1}{216}

OpenStudy (lolacole12):

got it and can i have a medal

OpenStudy (lolacole12):

@candice1997

OpenStudy (anonymous):

thank you for your help